OpenStudy (zuha):

Convert the map scale to a unit rate. Show your work. Interpret the meaning of the unit rate. 3/10 in = 7/8 mi

OpenStudy (anonymous):

3/10 = 7/8 1 - 10/10)= X 3/10x = 7/8 x= 7/8 divided by 3/10 = 7/8 x 10/3 = 70/24 = 2.916 And 1 unit = 2.916 units